# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 2.2.2 Unitary transformation of initial state
A related approach that was first suggested in [116], further developed in [117] and experimentally demonstrated on trapped ions in [9], is to accelerate the relaxation towards the equilibrium by a unitary transformation that sets $a_{2}=0$. The idea of first establishing $a_{2}=0$ and only then quenching the system to a thermal bath at $T_{\rm final}$ was discussed also in the classical setup in [65]; however, in that work, setting $a_{2}=0$ was done by coupling to a different bath (pre-heating or pre-cooling), whereas in the quantum scenario a natural approach is to use a unitary transformation instead. In principle, this technique is an application of Mpemba effect that belongs to Sec.(8) alongside with its classical analog. But since it played an important role in the study of quantum Mpemba effects, we review it here.
@@
The above quantity has the same sign as $\alpha_{1}$ for $s=0$ but, on the other hand, it has the same sign as $\alpha_{n}$ for $s=\pi/2$. In particular, it vanishes for $\bar{s}=\arctan\left(\sqrt{\left|\alpha_{1}/\alpha_{n}\right|}\right)$, so that if we take the unitary transformation $U=U(\bar{s})U_{1}$, Eq. (67) is satisfied. This implies that the initial state is rotated into a state that is orthogonal to the slowest decaying mode and will thus relax, in general, with the $1/|{\rm Re}(\lambda_{3})|$. In particular, this means that the approach to stationarity has been exponentially accelerated by a factor $|{\rm Re}(\lambda_{3})|-|{\rm Re}(\lambda_{2})|$.
Another way to obtain a quantum Markovian Mpemba effect is using a global unitary transformation when the previous transformation is not applicable. A very common case is when the eigenvalues of the lowest excited modes of the master operator form a complex conjugate pair. In this case, the initial state of the system approaches the steady state displaying oscillations in addition to exponential decay.
@@ 2.2.3 Entanglement asymmetry
A different type of effect that is also commonly referred to as the “quantum Mpemba effect” was recently suggested in [118], further studied in [119, 120, 121, 122, 123, 124, 125, 126] and experimentally demonstrated in a trapped ions setup [10]. In the terminology defined in Sec.  1.3, this effect can be categorized as a “non-thermal quantum Mpemba effect.” In this case, the dynamics of an isolated system rather than a system coupled to a bath is considered. In other words, the system is not coupled to any environment and follows Hamiltonian dynamics. To nevertheless observe some relaxation, the Hamiltonian system is divided onto two spatially separated parts. One part can be viewed as “the system”, whereas the other part serves as a “local bath,” but in contrast to the usual case the dynamics of this “bath” are also taken into account. Interestingly, this is closely related to the framework for classical macroscopic system used in [127] to study the Mpemba effect through a first-order phase transition, which we discussed in Sec. 3.1.
The following scenario is considered in this effect: The system, which includes both the “system” and the “bath,” is prepared in some initial condition, usually a pure state. The full system then evolves under a quantum Hamiltonian dynamics that has some global symmetry (e.g., $U(1)$ symmetry in [118, 10, 128], $S\!U(2)$ and $Z_{2}$ in [122], and even particle number symmetry [129] and “space-time symmetry” in [130]). The initial condition of the system is chosen such that the symmetry is not respected; namely, it is not an eigenstate of the system, which can be static or periodically driven Hamiltonian [131]. This asymmetry can be quantified by an “asymmetry parameter,” the entanglement entropy asymmetry, which is the relative entropy between the density matrix projected onto the conserved quantity subspace and the density matrix traced over the bath. During the dynamics, this asymmetry decays because of the global symmetry of the Hamiltonian, and thus the symmetry of the system is also restored. When the entanglement asymmetry decays faster for the initial condition that has the larger asymmetry in the initial condition, then there is a “quantum entanglement asymmetry Mpemba effect.” The terminology is chosen in analogy to the classical Mpemba effect. This scenario is demonstrated in the experimental measurements shown in Fig. 33 and discussed in Sec. 6.3.
@@ 2.3 Discussion
    Instantaneous Temperature Quench: In addition to the “perfect bath” assumption that implies Markovianity, our discussion assumed that the environment temperature is instantaneously changed from $T_{i}$ to $T_{f}$. This is a reasonable assumption in many cases, for both small and thermodynamically large systems, e.g., when suddenly placing a cup of water into a freezer; it is also a good description for the experimental system discussed in [6, 64, 8, 133]. However, this assumption is not always valid; in several experiments, the temperature changes from $T_{i}$ to $T_{f}$ at a (constant or varying) rate (e.g., 5 K/min in the differential scanning calorimetry used in [134]). When the temperature is not instantaneously quenched but rather changes gradually, the system can still be analyzed using a Markovian dynamics; however, in this case, the Markovian operator (Fokker-Planck operator or rate matrix) is not constant but rather changes with time. This complicates considerably the analysis, since the eigenvectors of the Markovian operator change, too. It is, therefore, possible that the coefficient along $\vec{v}_{2}$ corresponding to the final temperature operator is zero in the initial state, but evolution with a different operator that corresponds to a different temperature increases the corresponding coefficient. However, it is still possible to find the strong version of the Mpemba effects, using essentially the same idea presented in the example of colloidal setup with boundary coupling discussed in Sec. 2.1.7 (see Fig. 14). In this example, the operator is time-dependent because the temperature profile of the fluid varies with time.
@@
    Linearity. Assuming linear dynamics for the probability distribution is quite natural because Leuville’s equation for the probability distribution in Hamiltonian dynamics is a linear equation. Commonly, the dynamic of a system is modeled as a Hamiltonian dynamics plus dissipation and random forcing – both forces are generated as a result of the coupling of the system to the thermal environment. Assuming memory-less random force as well as dissipation, which is linear in the momentum, results in a linear dynamic for the probability distribution.
    The analysis presented in this section was heavily based on the eigenvectors and eigenvalues of the linear dynamics, therefore one might assume that the linear assumption is crucial for this analysis. However, this is not the case: even if the time evolution of the probability is nonlinear, we nevertheless assume that in the long time limit the probability relaxes to a unique steady state – the equilibrium distribution. From a dynamical system point of view, the equilibrium distribution can thus be considered as the unique stable fixed of the dynamic. It is, therefore, possible to linearize the dynamic in the long time limit and identify the slowest dynamics manifold from which all initial conditions relax towards the equilibrium point, except for the initial condition that relaxes from the codimension one “fast manifold” that has no coefficient along the slowest dynamic. Therefore, it is possible to identify the strong (direct and inverse) Mpemba effects even in this case. An example of nonlinear dynamics of the probability distribution was given in [96] for the mean-field anti-ferromagnetic with Glauber dynamics.
@@ 3 Phase transitions
Historically, the Mpemba effect was first discussed in the context of observations and experiments involving freezing water, and its initial definitions were often phrased in terms such as “the time to begin freezing.” However, the exact definition of the phase transition time in a first-order phase transition, where the two phases coexist for some nonzero fraction of the time, is tricky. Indeed, different definitions were used in different experiments to measure the “freezing time”: the moment at which the temperature at a specific point in the system (e.g., the point where the thermometer is placed) reaches the critical temperature [25] and the moment at which the latent heat begins to be released [22, 134]. Under some experimental conditions, the differences in the freezing time seem to be large enough that the exact definition does not play a role, and the effect exists regardless of the exact definition. However, for the case of water, the Mpemba effect is observed reliably only when freezing occurs and when its timing is a part of the definition [25, 50].
@@
From a theoretical perspective, defining the effect based on the phase transition time raises several difficulties, and indeed, the majority of works on the Mpemba effect used frameworks that are independent of a phase transition in the system. Some of the challenges that are raised by defining the effect based on a phase transition are: (i) First-order phase transitions are often dominated by rare events (e.g., in a non-homogeneous nucleation process). Consequently, there is a wide temporal distribution for the phase transition time. Estimating the phase transition time distribution when the process is dominated by a rare event is difficult theoretically, numerically and even experimentally. (ii) To observe the phase transition time, one needs to follow the dynamics of a large system that is far from the linear-response regime after a thermal quench. There are very few cases where analytical results for the dynamic are known, and these are often mean-field results. (iii) Even when it is possible to follow the dynamics of the system analytically or numerically, identifying the moment at which the phase transition happens is note easy. A possible way to bypass these issues is to use mean-field models. Indeed, all the works on Mpemba effects through a phase transition so far have used some kind of mean-field approximation and are thus quite far from being directly applicable to any of the experimental results.
@@
In what follows, we discuss separately anomalous relaxations through first- and second-order phase transitions. A key feature in first-order phase transitions is the existence of a metastable state. The transition from the metastable state to the global minimum of the free energy – the equilibrium state – is assumed to be the slowest process in the relaxation, and by analyzing this transition, it is possible to identify mechanisms for a Mpemba effect. The physical picture is quite different in second-order phase transitions, where there are no metastable states. A major obstacle, in this case, is to define a useful “phase transition time”; however, once such a time is determined, identifying a Mpemba effect is straightforward.
@@ 3.1 Mpemba effect through a first-order phase transition
Two different mechanisms were suggested to explain the Mpemba effect through a first-order phase transition. The first approach was suggested in [127] and further developed in [135]. It applies to any system with a metastable state associated with the first-order phase transition and coupled to a finite bath. The basic idea is as follows: When the system has a metastable state, the relaxation towards the final (stable) equilibrium state has two timescales: a fast relaxation of the system from its initial state to the metastable state, followed by a slow transition from the metastable state to the stable, equilibrium state. As intuitively expected, the initially hotter system reaches the metastable state faster (unless there is some type of a final relaxation Mpemba effect in this part of the relaxation, which, for simplicity, we assume is not the case). However, this time difference is much shorter than the timescale associated with the transition from the metastable state to the equilibrium state. The latter is a function of the state of the thermal bath, not only of the system. If the thermal bath is finite, it is affected by the relaxation of the system to the metastable state. Indeed, the initially hot system dumps more energy into the finite bath while cooling towards the metastable state; hence, the bath coupled to it has a higher effective temperature when the system reaches the metastable state than an identical thermal bath that was coupled to the same system that was initiated at a colder temperature. This difference in the bath temperature influences the transition rate from the metastable state to the stable equilibrium state: the hotter the bath, the faster the transition to the stable equilibrium state. Therefore, the initially hot system, which dumps more heat into the thermal bath during its relaxation to the metastable state, is coupled to an effectively higher temperature thermal bath; hence, its transition to the stable equilibrium state is faster.
The mechanism described above is quite generic. So far, it was demonstrated on a relatively simple model – the mean-field Ising model with a staggered magnetic field. The Hamiltonian of the system is given by
@@
where $\sigma_{i}\in\{-1,+1\}$ is a spin degree of freedom at site $i$, $K$ the staggered magnetic field magnitude, $J$ the two-spin coupling constant, and $I$ a four-spin coupling constant. Interestingly, this specific model is very similar to the antiferromagnet system studied in [5], which has a rich phase diagram of final relaxation Mpemba effects. Being a mean-field model, a microstate of the system can be uniquely characterized by two numbers as—for example, the staggered magnetization per spin $u=N^{-1}\sum_{i=1}^{N}(-1)^{i}\sigma_{i}$ and the total magnetization, $m=N^{-1}\sum_{i=1}^{N}\sigma_{i}$. The free energy landscape of this model with $K=-0.95$, $J=-1$, $I=1$ at $\beta=1.5385$ is shown in Fig. 16. It has a local minimum (the red point at $m=0$, $u=0.815$) that is a metastable state, whereas the global minima are at low/high values of $m$ (the black points in Fig. 16). During the first stage of the relaxation, the system stays in the metastable state for quite a long time until a rare fluctuation takes it to the final equilibrium state. The mean time for the system to go through the phase transition can be simply modeled as an Arrhenius law, $\tau\sim\exp(N\Delta E)$ [127], but a more sophisticated calculation [135] showed that it behaves as $\tau\sim\exp(\Delta S_{\textrm{tot}})$, where $\Delta S_{\textrm{tot}}$ is the total entropy change (system + bath) between the metastable state and the global minimum.
Figure 16: Level sets of the free energy, $F(u,m)$, for the model in Eq. (73), as a function of staggered magnetization per spin, $u$, and magnetization per spin, $m$, at $K=-0.95$, $J=-1$, $I=1$ and $\beta=1.5385$. The red point is a local minimum of $F(u,m)$, which is a metastable state. The global minima are the two black points. The green line is the equilibrium line of the system at these values of $K$, $J$, and $I$ – it has a first-order phase transition, which is the abrupt jump in the equilibrium line from $m=0$ to $m\neq 0$. A system initiated at the hot phase with $m=0$ first relaxes quickly towards the metastable red point and then more slowly to equilibrium (black points).
@@
To demonstrate the effect in this model, a microcanonical Monte Carlo simulation (the Creutz algorithm) for a system of $N=400$ spins in [127] and $N=50$ spins in [135], coupled to a finite bath of with $N_{\textrm{bath}}=20N$ in [127] and $N_{\textrm{bath}}=50N$ in [135] was made. The system was initiated at the equilibrium distributions for various initial temperatures, and the bath was sampled from a temperature far below the phase transition temperature. A phase transition was declared when the mean magnetization “jumped” from the local, metastable minimum that has zero magnetization to one of the two global minima of non-zero magnetization (see Fig. 17). The numerical demonstrations show that, indeed, the initially hot system goes through the phase transition before the initially cold system.
@@
Figure 17: Creutz microcanonical Monte Carlo simulations of the system described by Eq. (73). The black and gray graphs show the mean magnetization per spin (in absolute value) as a function of the Monte Carlo steps for two initial conditions sampled from different temperatures. The phase transition is declared by the abrupt jump from fluctuations around $|m|=0$ to $|m|\sim 1$. The figure from reproduced from [127].
This mechanism for a Mpemba effect through a first-order phase transition is quite generic: first-order phase transitions are commonly associated with a metastable state, and many such systems coupled to a finite bath should show a similar effect, since generically, the initially hotter system dumps more energy to the thermal bath and therefore is coupled to a bath with a higher temperature which then makes the escape from the metastable state faster. Clearly, this approach is not relevant for an inverse Mpemba effect (which so far was never measured through a first-order phase transition), since in such a scenario the cold initial condition gets more heat out of the finite thermal bath; therefore, when it reaches the metastable state it is coupled to a colder bath, which makes the escape slower.
@@
There are a few difficulties with this approach: (1) The phase transition time, as well as its difference between the cold and hot initial conditions, diverges exponentially in the thermodynamic limit, but this limit is essential for the existence of a phase transition. (2) The existence of the effect strongly depends on the finite size of the bath, even though some effect should exist even when the thermal bath is large [135]. (3) This mechanism is useful for mean-field models, where the phase transition is not dominated by nucleation (homogeneous or non-homogeneous) processes, by the coupling of the system to the bath, and other factors that are known to be important in practice. In finite-dimension systems, the relaxation from the metastable state to the final equilibrium might be affected by additional factors.
@@
A carefully constructed mean field Blume-Emery-Griffiths model was used to demonstrate this effect. Because the model is mean field, the state of the system in the thermodynamic limit is fully characterized by two macroscopic parameters ($x_{1}$ and $x_{2}$ in Fig. 18). The parameters of the system were thus chosen such that the hot initial condition is on the ridge between the basin of attractions corresponding to the two minima, while the medium temperature initial condition is in the basin of attraction of the metastable state (see Fig. 18a). Thus, the time to escape the unstable initial condition of the hot system grows with the logarithm of the system size, whereas the time to exist in the metastable state in the medium temperature initial condition grows exponentially with system size.
@@
Figure 18: The free energy, $F(x_{1},x_{2})$, landscape of the Blume-Emery-Griffiths mean-field model. (a) $F(x_{1},x_{2})$ for $T=T_{\mathrm{final}}$. This landscape has three minima: two symmetric global minima $\vec{x}(T_{f})$ (brown points) corresponding to the equilibrium state at $T_{\mathrm{final}}$, and one local minimum (green point). Each minimum has a basin of attraction, and the three basins are separated by ridges. The initial condition at the hot temperature $\vec{x}(T_{h})$ (red point) is on the ridge between the basins of the two equilibrium points, whereas the cold initial condition, $\vec{x}(T_{c})$ (blue point) is in the basin of attraction of the metastable local minimum. Therefore, it will take the cold initial condition much longer time to relax towards the equilibrium state of the model. Panel (b): The equilibrium line of the model at the relevant parameters in the $(x_{1},x_{2})$ plane. Panel (c): Free energy $F(x_{1},x_{2})$ at temperature $T_{h}$. These figures are reproduced from [58].
@@
This mechanism is quite generic and might exist in any system that has the relevant structure of metastable state and initial equilibrium distributions that flow into the equilibrium or metastable state basins of attraction. It bypasses the need to carefully define the phase transition time in a way that does not diverge in the thermodynamic limit since, regardless of how it is defined, it is clear that initiating the system in the basin of attraction of a metastable state for relatively cold initial conditions and of the equilibrium state for hot initial conditions can generate the effect. On the other hand, it is also clear that this effect has drastically different characteristics from the effects observed experimentally: for this mechanism to exist, there should be a clear range of parameters at which the system relaxes to a metastable state, and a range of parameters from which it never does. It is very probable that this mechanism exists in some experimental systems; however, it is challenging to analyze it theoretically or even numerically beyond a mean-field model, namely in a model whose state during relaxation depends on many macroscopic order parameters. In addition, we lack a simple physical picture to connect the basin of attractions of the metastable and equilibrium states to the hot/medium temperature initial conditions. Such a connection between the equilibrium line and the basin of attraction at low temperature – which is a property of the dynamics, not just the equilibrium state of the model – might also shed light on how generic the effect is: Is it expected in many models, or does it require some careful tuning or special mechanism?

### v3

@@ 2.2.2 Unitary transformation of initial state
A related approach that was first suggested in [121], further developed in [122] and experimentally demonstrated on trapped ions in [9], is to accelerate the relaxation towards the equilibrium by a unitary transformation that sets $a_{2}=0$. The idea of first establishing $a_{2}=0$ and only then quenching the system to a thermal bath at $T_{\rm final}$ was discussed also in the classical setup in [65]; however, in that work, setting $a_{2}=0$ was done by coupling to a different bath (pre-heating or pre-cooling), whereas in the quantum scenario a natural approach is to use a unitary transformation instead. In principle, this technique is an application of Mpemba effect that belongs to Sec. 8 along with its classical analog. But since it played an important role in the study of quantum Mpemba effects, we review it here.
@@
The above quantity has the same sign as $\alpha_{1}$ for $s=0$ but has the same sign as $\alpha_{n}$ for $s=\pi/2$. In particular, it vanishes for $\bar{s}=\arctan\left(\sqrt{\left|\alpha_{1}/\alpha_{n}\right|}\right)$, so that if we take the unitary transformation $U=U(\bar{s})U_{1}$, Eq. (67) is satisfied. This implies that the initial state is rotated into one that is orthogonal to the slowest decaying mode and will thus relax, in general, on a timescale $1/|{\rm Re}(\lambda_{3})|$. The approach to stationarity has thus been exponentially accelerated by a factor $|{\rm Re}(\lambda_{3})|-|{\rm Re}(\lambda_{2})|$.
Another way to obtain a quantum Markovian Mpemba effect is using a global unitary transformation when the previous transformation is not applicable. A very common case is when the eigenvalues of the lowest excited modes of the master operator form a complex-conjugate pair. In this case, the evolving system state oscillates as it decays to steady state.
@@ 2.2.3 Entanglement asymmetry
A different type of effect that is also commonly referred to as the “quantum Mpemba effect” was recently suggested in [123], further studied in [124, 125, 126, 127, 128, 129, 130, 131] and experimentally demonstrated in a trapped ions setup [10]. In the terminology defined in Sec. 1.3, this effect can be categorized as a “non-thermal quantum Mpemba effect.” In this case, the dynamics of an isolated system rather than a system coupled to a bath is considered. In other words, the system is not coupled to any environment and follows Hamiltonian dynamics. To nevertheless observe some relaxation, one divides the Hamiltonian system into two spatially separated parts. One part can be viewed as “the system,” whereas the other part serves as a “local bath”; however, in contrast to the usual case, the dynamics of this “bath” are also taken into account. Interestingly, this is closely related to the framework for classical macroscopic system used in [132] to study the Mpemba effect through a first-order phase transition, which we discussed in Sec. 3.1.
The following scenario is considered in this effect: The system, which includes both the “system” and the “bath,” is prepared in some initial condition, usually a pure state. The full system then evolves under a quantum Hamiltonian dynamics that has some global symmetry. For example, the particle number symmetry, i.e. $U(1)$, in [123, 10, 133, 134, 135], $S\!U(2)$ and $Z_{2}$ in [136, 127] and “space-time symmetry” in [137]. The initial condition of the system is chosen such that the symmetry is not respected; namely, it is not an eigenstate of the system, which can be static or periodically driven Hamiltonian [138]. This asymmetry can be quantified by an “asymmetry parameter,” the entanglement asymmetry, which is the relative entropy between the density matrix projected onto the conserved quantity subspace and the density matrix traced over the bath. During the dynamics, this asymmetry decays because of the global symmetry of the Hamiltonian, and thus the symmetry of the system is also restored. When the entanglement asymmetry decays faster for the initial condition with the larger asymmetry in the initial condition, there is a “quantum entanglement asymmetry Mpemba effect.” The terminology is chosen in analogy to the classical Mpemba effect. This scenario is demonstrated in the experimental measurements shown in Fig. 33 and discussed in Sec. 6.3.
@@ 2.3 Discussion
    Instantaneous Temperature Quench: In addition to the “perfect bath” assumption that implies Markovianity, our discussion assumed that the environment temperature is instantaneously changed from $T_{i}$ to $T_{f}$. This is a reasonable assumption in many cases, for both small and thermodynamically large systems, e.g., when suddenly placing a cup of water into a freezer; it is also a good description for the experimental system discussed in [6, 64, 8, 140]. However, this assumption is not always valid; in several experiments, the temperature changes from $T_{i}$ to $T_{f}$ at a (constant or varying) rate (e.g., 5 K/min in the differential scanning calorimetry used in [141]). When the temperature is not instantaneously quenched but rather changes gradually, the system can still be analyzed using a Markovian dynamics; however, in this case, the Markovian operator (Fokker-Planck operator or rate matrix) is not constant but rather changes with time. This complicates considerably the analysis, since the eigenvectors of the Markovian operator change, too. It is therefore possible that the coefficient along $\vec{v}_{2}$ corresponding to the final temperature operator is zero in the initial state, but evolution with a different operator that corresponds to a different temperature increases the corresponding coefficient. However, it is still possible to find the strong version of the Mpemba effects, using essentially the same idea presented in the example of colloidal setup with boundary coupling discussed in Sec. 2.1.7; see Fig. 14. In this example, the operator is time dependent because the temperature profile of the fluid varies with time.
@@
    Linearity. Assuming linear dynamics for the probability distribution is quite natural because Liouville’s equation for the probability distribution in Hamiltonian dynamics is linear. Commonly, the dynamics of a system are modeled as Hamiltonian dynamics plus dissipation and random forcing; both forces are generated as a result of the coupling of the system to the thermal environment. Assuming memory-less random force as well as dissipation, which is linear in the momentum, results in linear dynamics for the probability distribution.
    Since the analysis presented in this section was heavily based on the eigenvectors and eigenvalues of the linear dynamics, one might worry that the linear assumption is crucial for the analysis. However, this is not the case: even if the time evolution of the probability is nonlinear, we nevertheless assume that in the long-time limit the probability relaxes to a unique steady state, the equilibrium distribution. From a dynamical-system point of view, the equilibrium distribution can thus be considered as the unique stable fixed point of the dynamics. It is, therefore, possible to linearize the dynamics in the long-time limit and identify the slowest-dynamics manifold from which all initial conditions relax towards the equilibrium point, except for the initial condition that relaxes from the codimension-one “fast manifold” that has no coefficient along the slowest dynamical variable. Therefore, it is possible to identify strong direct and inverse Mpemba effects, even in this case. An example of nonlinear dynamics of the probability distribution was given in [96] for the mean-field anti-ferromagnetic with Glauber dynamics.
@@ 3 Phase transitions
Historically, the Mpemba effect was first discussed in the context of observations and experiments involving freezing water, and its initial definitions were often phrased in terms such as “the time to begin freezing.” However, the exact definition of the phase transition time in a first-order phase transition, where the two phases coexist for some nonzero fraction of the time, is tricky. Indeed, different definitions were used in different experiments to measure the “freezing time,” one being the moment at which the temperature at a specific point in the system (e.g., the point where the thermometer is placed) reaches the critical temperature [25] and another being the moment at which the latent heat begins to be released [22, 141]. Under some experimental conditions, the differences in freezing time seem large enough that the exact definition does not play a role, and the effect exists regardless of the exact definition. However, for the case of water, the Mpemba effect is observed reliably only when freezing occurs and when its timing is a part of the definition [25, 50].
@@
From a theoretical perspective, defining the effect based on the phase transition time raises several difficulties, and indeed, the majority of works on the Mpemba effect used frameworks that are independent of a phase transition in the system. Some of the challenges raised by incorporating phase transitions into the definition of the Mpemba effect are as follows: (i) First-order phase transitions are often dominated by rare events (especially in an inhomogeneous nucleation process). Consequently, there is a wide temporal distribution for the phase-transition time. Estimating the phase-transition-time distribution when the process is dominated by a rare event is difficult theoretically, numerically, and even experimentally. (ii) To observe the phase transition time, one needs to follow the dynamics of a large system that is far from the linear-response regime after a thermal quench. There are very few cases for which analytical results for the dynamic are known, and these are often mean-field results. (iii) Even when it is possible to follow the dynamics of the system analytically or numerically, identifying the moment at which the phase transition happens is not easy. A possible way to bypass these issues is to use mean-field models. Indeed, all the works on Mpemba effects through a phase transition to date have used some kind of mean-field approximation and are thus quite far from being directly applicable to any of the experimental results.
@@
In what follows, we discuss separately anomalous relaxations through first- and second-order phase transitions. A key feature in first-order phase transitions is the existence of a metastable state. The transition from the metastable state to the global minimum of the free energy—the equilibrium state—is assumed to be the slowest process in the relaxation, and by analyzing this transition, it is possible to identify mechanisms for a Mpemba effect. The physical picture differs in second-order phase transitions, where there are no metastable states. A major obstacle, in this case, is to define a useful “phase transition time”; however, once such a time is determined, identifying a Mpemba effect is straightforward.
@@ 3.1 Mpemba effect through a first-order phase transition
Figure 16: Level sets of the free energy $F(u,m)$ for the mean-field Ising model described by Eq. (73), as a function of staggered magnetization per spin $u$ and magnetization per spin $m$, at $K=-0.95$, $J=-1$, $I=1$ and $\beta=1.5385$. The red marker indicates a local minimum of $F(u,m)$, which is a metastable state. The global minima are denoted by the two black markers. The two green lines represent the equilibrium locus of the system for these values of $K$, $J$, and $I$. As the system has a first-order phase transition, there is an abrupt jump from $m=0$ to $m\neq 0$, which splits the equilibrium locus to two disjoint lines. A system initiated at the hot phase with $m=0$ first relaxes quickly towards the metastable state (the red marker) and then more slowly to equilibrium (the black markers).
Two different mechanisms have been proposed to explain the Mpemba effect through a first-order phase transition. The first, suggested in [132] and developed in [142], applies to any system with a metastable state associated with the first-order phase transition and coupled to a finite bath. Recall from Sec. 2.1.8 that a system with a metastable state relaxes following two timescales towards the final, stable equilibrium state: A fast relaxation (i) of the system from its initial state to the metastable state is followed by a slow transition (ii) from the metastable state to the stable, equilibrium state. The first stage, whether starting from a hot or cold state, is typically much faster than the second stage. But—and this is the key point—if the thermal bath is finite, then the energy transferred from system to bath during the first stage will increase the bath temperature. The initially hot system, which transfers more heat into the thermal bath during its relaxation to the metastable state, is thus coupled to an effectively higher temperature thermal bath and hence reaches sooner the stable equilibrium state.
The mechanism described above is generic. So far, it has been demonstrated on a relatively simple model, the mean-field Ising model with staggered magnetic field; see Fig. 16. The Hamiltonian of the system is given by
@@
where $\sigma_{i}\in\{-1,+1\}$ is a spin degree of freedom at site $i$, $K$ the staggered magnetic field magnitude, $J$ the two-spin coupling constant, and $I$ a four-spin coupling constant. Interestingly, this specific model is very similar to the antiferromagnet system studied in [5], which has a rich phase diagram of final relaxation Mpemba effects. Being a mean-field model, a microstate of the system can be uniquely characterized by two numbers, for example the staggered magnetization per spin $u=N^{-1}\sum_{i=1}^{N}(-1)^{i}\sigma_{i}$ and the total magnetization, $m=N^{-1}\sum_{i=1}^{N}\sigma_{i}$. The free energy landscape of this model with $K=-0.95$, $J=-1$, $I=1$ at $\beta=1.5385$ is shown in Fig. 16. It has a local minimum (the red marker at $m=0$, $u=0.815$) that is a metastable state, whereas the global minima are at low/high values of $m$ (the black markers at top and bottom of Fig. 16). During the first stage of the relaxation, the system stays in the metastable state for a long time until a rare fluctuation takes it to the final equilibrium state. The mean time for the system to go through the phase transition can be simply modeled as an Arrhenius law, $\tau\sim\exp(N\Delta E)$ [132], but a more sophisticated calculation [142] showed that $\tau\sim\exp(\Delta S_{\textrm{tot}})$, where $\Delta S_{\textrm{tot}}$ is the total entropy change (for system and bath) between the metastable state and the global minimum.
@@
The effect in this model was demonstrated via a microcanonical Monte Carlo simulation, based on the Creutz algorithm, for a system of $N=400$ spins in [132] and $N=50$ spins in [142]. The spins were coupled to a finite bath, with $N_{\textrm{bath}}=20N$ in [132] and $50N$ in [142]. The system was initiated at the equilibrium distributions for various initial temperatures, and the bath was sampled from a temperature far below the phase transition temperature. A phase transition was identified by the “jump” in mean magnetization from the local, metastable minimum at zero magnetization to one of the two global minima at non-zero magnetization; see Fig. 17. Indeed, the initially hot system is observed to go through the phase transition before the initially cold system.
@@
Figure 17: Creutz microcanonical Monte Carlo simulations of the system described by Eq. (73). The black and gray graphs show the mean magnetization per spin, in absolute value, as a function of the Monte Carlo steps for two initial conditions sampled from different temperatures. The phase transition is declared by the abrupt jump from fluctuations around $|m|=0$ to $|m|\sim 1$. Source: Reprinted with permission from [132].
This mechanism for a Mpemba effect through a first-order phase transition is generic: first-order phase transitions are usually associated with a metastable state, and finite baths are common. Clearly, however, this approach is not relevant for an inverse Mpemba effect, where the signs are all reversed. That is, the colder initial state transfers more heat out of the thermal bath, reducing its temperature and thereby slowing the second stage. We note that an inverse Mpemba effect has never been measured through a first-order phase transition.
@@
There are also a few difficulties with this proposed mechanism: (1) The phase transition time, as well as its difference between the cold and hot initial conditions, diverges exponentially in the thermodynamic limit, but this limit is essential for the existence of a phase transition. (2) The existence of the effect strongly depends on the finite size of the bath, even though some effect should exist even when the thermal bath is large [142]. (3) This mechanism is useful for mean-field models, where the phase transition is not dominated by nucleation (homogeneous or inhomogeneous) processes, by the coupling of the system to the bath, and other factors that are known to be important in practice. In finite-dimension systems, the relaxation from the metastable state to the final equilibrium might be affected by additional factors.
@@
A carefully constructed mean field Blume-Emery-Griffiths model was used to demonstrate this effect. Because the model is mean field, the state of the system in the thermodynamic limit is fully characterized by two macroscopic parameters ($x_{1}$ and $x_{2}$ in Fig. 18). The parameters of the system were thus chosen such that the hot initial condition is on the ridge between the basin of attractions corresponding to the two minima, while the medium temperature initial condition is in the basin of attraction of the metastable state; see Fig. 18a. Thus, the time to escape the unstable initial condition of the hot system grows with the logarithm of the system size, whereas the time to exist in the metastable state in the medium temperature initial condition grows exponentially with system size.
@@
Figure 18: The free energy, $F(x_{1},x_{2})$, landscape of the Blume-Emery-Griffiths mean-field model. (a) $F(x_{1},x_{2})$ for $T=T_{\mathrm{final}}$. This landscape has three minima: two symmetric global minima $\vec{x}(T_{f})$ (brown points) corresponding to the equilibrium state at $T_{\mathrm{final}}$, and one local minimum (green point). Each minimum has a basin of attraction, and the three basins are separated by ridges. The initial condition at the hot temperature $\vec{x}(T_{h})$ (red point) is on the ridge between the basins of the two equilibrium points, whereas the cold initial condition, $\vec{x}(T_{c})$ (blue point) is in the basin of attraction of the metastable local minimum. Therefore, it will take the cold initial condition much longer time to relax towards the equilibrium state of the model. Panel (b): The equilibrium line of the model at the relevant parameters in the $(x_{1},x_{2})$ plane. Panel (c): Free energy $F(x_{1},x_{2})$ at temperature $T_{h}$. Source: Reprinted with permission from [58].
@@
This mechanism is generic and might exist in any system that has the relevant structure of metastable state and initial equilibrium distributions that flow into the equilibrium or metastable state basins of attraction. It bypasses the need to carefully define the phase transition time in a way that does not diverge in the thermodynamic limit since, regardless of how it is defined, it is clear that initiating the system in the basin of attraction of a metastable state for relatively cold initial conditions and of the equilibrium state for hot initial conditions can generate the effect. On the other hand, it is also clear that this effect has drastically different characteristics from the effects observed experimentally: for this mechanism to exist, there should be a clear range of parameters at which the system relaxes to a metastable state, and a range of parameters from which it never does. It seems likely that this mechanism exists in some experimental systems; however, it is challenging to analyze it theoretically or even numerically beyond a mean-field model, namely in a model whose state during relaxation depends on many macroscopic order parameters. In addition, we lack a simple physical picture to connect the basin of attractions of the metastable and equilibrium states to the hot/cold temperature initial conditions. Such a connection between the equilibrium line and the basin of attraction at low temperatures, which is a property of the dynamics and not just the equilibrium state of the model, might also shed light on how generic is the effect: Is it expected in many models, or does it require some careful tuning or special mechanism?